OpenStudy (anonymous):

Solve the equation for x and round your answer to the nearest hundredth. 5^x + 14 = 26

OpenStudy (anonymous):

14 and 26

OpenStudy (anonymous):

The answer has to be a decimal i figured it out i think it's 1.54

OpenStudy (yttrium):

5^x = 12 log base (5) 12= x x = log12/log5 x = log (4*3)/log5 x = log(2^2 * 3)/log5 x = (2log2 + log3)/log5 Is this what you did?